1. The half-life of an element is 30 years. We start with 50 grams. Find a formula that gives the amount of the element remaining after h half-lives.

Select one: a. 50 cross times 0.5 to the power of h

b. 50 minus 0.5 cross times h

c. 15 to the power of h

d. 15 plus 0.5 to the power of h

e. None of these

2.

A bank account starts with $ 10. Every month, the amount in the account is tripled. Give a formula of the balance after t months.

Select one:

a.

b.

c.

d. None of these

e.

3.

How many times more intense is a 5.0-magnitude earthquake than a 3.0-magnitude
earthquake?

Select one:

a. 100

b. 2

c. None of these

d. 1000

e. 3

4.

Suppose you invest $5000 in a savings account that pays an APR of 4%. If the interest is
compounded monthly, what is the balance in the account after 10 years?

Select one:

a. None of these

b. $5169.19

c. $8072.64

d. $7454.16

e. $5203.71

5.

Given a set of data points, the _____ line comes as close as possible to fitting those data.

Select one:

a. growth

b. slope

c. None of these

d. interpretation

e. regression
Answer #1

Answer (1) ACH) = Ao ekt Ao : 50 -kt so A(t) = 50e 25 A(30) : 40 -30K 25 = 50e e 30k 0.5 . - 30K in (0.5): ince 30%) In (0.5)log (2/3) = logx-logy 109풍 - 109를 :5-3 (109 I, -1095) - (log I, - 1095): 2 109 I, - 109 I 2 = 2 log ( 7 ) I Iz - 10 (cogeny x